OpenStudy (anonymous):

How would you convert the repeating, nonterminating decimal to a fraction? Explain the process as you solve the problem 0.1515...

OpenStudy (danjs):

The way i remember, is to use multiples of the thing like this.....

OpenStudy (danjs):

Let S = 0.151515.... so, 100*S = 15.151515... Subtract the two, the repeating decimal is gone, 100S - S = 15.151515... - 0.151515... 99S = 15 S = 15/99

OpenStudy (danjs):

you have to use different multiples of the decimal depending on the number, so that you can get to subtracting the repeating decimal from itself